You have been asked to plate cells in 50uL total volume on day 1 in every well of a 96 well plate. The second day, you will need to add 50uL of media containing drug so that your final volume in each well is 100 uL. Per well, you want 5000 cells and the drug at 1 uM final concentration. You are given 5 mL of cells at 1.3x10^7 cells/mL and 100 uL of stock drug solution at 10 mM. Please describe how you would make cell suspensions and media containing drug such that you have enough for all 96 wells. Be specific about the volumes you would use to dilute both the cell suspension and the drug solution. Assume you have a large amount of fresh media.
Answer #1

Given that we need to plate 5000 cells per well and a total of 96 wells we need to plate.

So the number of cells in 96 well plates will be 96*5000 = 480000 cells.

Volume to be plated in each well is 50 ul. So total volume = 96 * 50 = 4800ul or 4.8ml.

So we need 480000 cells in 4.8ml culture or 100000 cells/ml or 105 cells/ml

Cells present in original culture is 1.3*107cells/ml.

So we need to dilute it 1.3*107/105 = 130 times

or 1 ul of cell culture and 129 ul of DMEM.

We need 37 ul of cell culture and 4763 ml DMEM to make 5000 cells/ ul

For drug -

We need the final concentration of drug to be 1uM

Stock concentration is 10mM.

Given that we need to make the final concentration of drug for 100ul be 1uM.

total volume for 96 well plate will be 100*96 = 9.6ml or 9600ul.

C1V1 = C2V2

C1 = stock concentration of the drug = 10mM or 10000uM

V1 = volume needed to make final concentration of 1uM

C2 = 1uM

V2 = 9600 ul

10000 * V1 = 1 * 9600

V1 = 9600/10000 = 0.96ul.

Since 96-well plate has 50ul of DMEM in per well. So we will add extra 50ul to make total volume 100ul.

0.96ul of the drug from the stock solution will be added to the 4800ul of DMEM.

Please write to me if you need more help.